OpenStudy (anonymous):

Estimate the square root to the nearest integer -(squareroot)51

OpenStudy (anonymous):

what squares/sqrts do you know close to 51?

OpenStudy (unklerhaukus):

\[\sqrt{49}<\sqrt{51}<\sqrt{64}\]